# Defining subsystems in Hilbert spaces with non-Euclidean metric

### v1

@@ Appendix C Ambiguities in entropy of the state
Now, if the state $\omega$ is pure, so is the density matrix $\rho_{\omega}$. If it is a mixed state, it can be written as a convex combination of other pure states: $\omega=\sum_{\alpha}\lambda_{\alpha}\omega^{\alpha}$, and the corresponding density matrix can be written as $\rho_{\omega}=\sum_{\alpha}\lambda_{\alpha}{\rho_{\omega}}^{\alpha}$ where $\rho_{\omega}^{\alpha}$. are pure state density matrices Furthermore, the representation is reducible into a direct sum of irreducible representations $\rho_{\omega}=\oplus_{i}p_{i}\rho_{i}$. Note that here $\rho_{i}$ need not be pure states. The convex sum decomposition of the state $\omega$ is not unique since the state space is not a simplex. Therefore, a state can be prepared from an ensemble of pure states in many ways, represented by the probability vector $\vec{\lambda}$ s.t. $\sum_{\alpha}\lambda_{\alpha}=1$. The entropy associated with such a preparation is given by the Shannon entropy of the probability vector $\vec{\lambda}$: $H(\vec{\lambda})=-\sum_{\alpha}\lambda_{\alpha}\ln(\lambda_{\alpha})$ The von-Newmann entropy of the density matrix $\rho_{\omega}$, on the other hand, is given by $S(\rho_{\omega})=-tr(\rho_{\omega}\ln\rho_{\omega})=\sum_{i}e_{i}\ln e_{i}$ where $\vec{e_{i}}$ is the spectrum of the state $\rho_{\omega}$. The ambiguity in the entropy of the preparation of the state $\omega$ is related to the fact that a density matrix can be written as a convex combination of pure density matrices in many non-unique ways: $\rho=\sum_{i}\lambda_{i}|\psi_{i}\rangle\langle\psi_{i}|$. It turns out that the Shannon entropy of all such preparations (given by $H(\vec{\lambda})$ ) is bounded below by the von-Newmann entropy of the state: $S(\rho)=\inf\{H(\vec{\lambda});\rho=\lambda_{i}|\psi_{i}\rangle\langle\psi_{i}|\}$. Similarly one can construct the unique entropy of the state $\omega$ on the algebra $\mathcal{A}$ as the infimum of all the preparations of $\omega$: $S(\omega)=\inf\{H(\vec{\lambda});\omega=\sum_{\alpha}\lambda_{\alpha}\omega^{\alpha}\}$. We therefore have a unique notion of the entropy of a state if $S(\omega)=S(\rho_{\omega})$. As shown in [33], this is true only under faithful representations with no multiplicities. The most general relation between the two entropies is given as follows:
@@
Appendix D Spin tomography
General spin tomography: Given eigenvectors $|m\rangle$ of $\sigma_{z}$ with eigenvalues $m\in\{\dfrac{1}{2},-\dfrac{1}{2}\}$, the probability of obtaining the result $m$ after measuring the spin $\sigma_{z}$ is given by $p(m|\sigma_{z})=\langle m|\rho|m\rangle$. Consider a unitary evolution of this state under $U=(\mathbb{1}-i\sigma_{Y})/\sqrt{2}$, the subsequent measurement of spin in the $\sigma_{3}$ direction, the probability of getting the result $m$, given by $p(m)=\langle m|U^{\dagger}\rho U|m\rangle$ is the probability of getting the result $m$ after a $\sigma_{1}$ measurement on $\rho$. This is because $U|m\rangle$ is the eigenvector of $\sigma_{x}$ with eigenvalue $m$. In general, if one wants to measure the observable $\vec{\sigma}.\vec{n}$ where $\vec{n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, it is equivalent to a $\sigma_{z}$ measurement of the state $U\rho U^{\dagger}$ where $U=e^{-i(\theta/2)\sigma.\hat{n}_{\perp}}$ is the unitary with $\hat{n}_{\perp}=(-\sin\phi,\cos\phi,0)$. This is possible due to the following identity: $e^{-i(\theta/2)\sigma.\hat{n}_{\perp}}\sigma_{z}e^{i(\theta/2)\sigma.\hat{n}_{\perp}}=\vec{\sigma}.\vec{n}$  . In the Stern-Gerlach setting, the interaction Hamiltonian is given by $H=-\vec{\mu}.\vec{B}$, where the $\vec{\mu}=\gamma\vec{\sigma}$ is the magnetic moment and $\gamma$ is the gyromagnetic factor, and $\vec{B}$ is the magnetic field. The evolution $U$ is carried out by fixing the this evolution is carried out by the magnetic field $\vec{B}=\theta\hat{n}_{\perp}/(\gamma t)$.

### v2

@@ Appendix B Ambiguities in entropy of the state
Now, if the state $\omega$ is pure, so is the density matrix $\rho_{\omega}$. If it is a mixed state, it can be written as a convex combination of other pure states: $\omega=\sum_{\alpha}\lambda_{\alpha}\omega^{\alpha}$, and the corresponding density matrix can be written as $\rho_{\omega}=\sum_{\alpha}\lambda_{\alpha}{\rho_{\omega}}^{\alpha}$ where $\rho_{\omega}^{\alpha}$. are pure state density matrices Furthermore, the representation is reducible into a direct sum of irreducible representations $\rho_{\omega}=\oplus_{i}p_{i}\rho_{i}$. Note that here $\rho_{i}$ need not be pure states. The convex sum decomposition of the state $\omega$ is not unique since the state space is not a simplex. Therefore, a state can be prepared from an ensemble of pure states in many ways, represented by the probability vector $\vec{\lambda}$ s.t. $\sum_{\alpha}\lambda_{\alpha}=1$. The entropy associated with such a preparation is given by the Shannon entropy of the probability vector $\vec{\lambda}$: $H(\vec{\lambda})=-\sum_{\alpha}\lambda_{\alpha}\ln(\lambda_{\alpha})$ The von-Newmann entropy of the density matrix $\rho_{\omega}$, on the other hand, is given by $S(\rho_{\omega})=-tr(\rho_{\omega}\ln\rho_{\omega})=\sum_{i}e_{i}\ln e_{i}$ where $\vec{e}$ is the spectrum of the state $\rho_{\omega}$. A density matrix can be written as a convex combination of pure density matrices in many non-unique ways: $\rho=\sum_{i}\lambda_{i}|\psi_{i}\rangle\langle\psi_{i}|$. It turns out that the Shannon entropy of all such preparations (given by $H(\vec{\lambda})$ ) is bounded below by the von-Newmann entropy of the state: $S(\rho)=\inf\{H(\vec{\lambda});\rho=\lambda_{i}|\psi_{i}\rangle\langle\psi_{i}|\}$. Similarly one can construct the unique entropy of the state $\omega$ on the algebra $\mathcal{A}$ as the infimum of all the preparations of $\omega$: $S(\omega)=\inf\{H(\vec{\lambda});\omega=\sum_{\alpha}\lambda_{\alpha}\omega^{\alpha}\}$. We therefore have a unique notion of the entropy of a state if $S(\omega)=S(\rho_{\omega})$. As shown in [34], this is true only under faithful representations with no multiplicities. The most general relation between the two entropies is given as follows:
@@
Appendix C Spin tomography
General spin tomography: Given eigenvectors $|m\rangle$ of $\sigma_{z}$ with eigenvalues $m\in\{\dfrac{1}{2},-\dfrac{1}{2}\}$, the probability of obtaining the result $m$ after measuring the spin $\sigma_{z}$ is given by $p(m|\sigma_{z})=\langle m|\rho|m\rangle$. Consider a unitary evolution of this state under $U=(\mathbb{1}-i\sigma_{Y})/\sqrt{2}$, the subsequent measurement of spin in the $\sigma_{3}$ direction, the probability of getting the result $m$, given by $p(m)=\langle m|U^{\dagger}\rho U|m\rangle$ is the probability of getting the result $m$ after a $\sigma_{1}$ measurement on $\rho$. This is because $U|m\rangle$ is the eigenvector of $\sigma_{x}$ with eigenvalue $m$. In general, if one wants to measure the observable $\vec{\sigma}.\vec{n}$ where $\vec{n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, it is equivalent to a $\sigma_{z}$ measurement of the state $U\rho U^{\dagger}$ where $U=e^{-i(\theta/2)\sigma.\hat{n}_{\perp}}$ is unitary with $\hat{n}_{\perp}=(-\sin\phi,\cos\phi,0)$. This is possible due to the following identity: $e^{-i(\theta/2)\sigma.\hat{n}_{\perp}}\sigma_{z}e^{i(\theta/2)\sigma.\hat{n}_{\perp}}=\vec{\sigma}.\vec{n}$  . In the Stern-Gerlach setting, the interaction Hamiltonian is given by $H=-\vec{\mu}.\vec{B}$, where $\vec{\mu}=\gamma\vec{\sigma}$ is the magnetic moment, $\gamma$ is the gyromagnetic factor, and $\vec{B}$ is the magnetic field. The evolution $U$ is carried out by fixing the this evolution is carried out by the magnetic field $\vec{B}=\theta\hat{n}_{\perp}/(\gamma t)$.